OpenStudy (anonymous):

ln(8x-7)/ln(x)=2 please help!!

OpenStudy (ranga):

ln(8x-7) / ln(x) = 2 multiply both sides by ln(x) ln(8x - 7) = 2 * ln(x) = ln(x^2) Therefore, 8x - 7 = x^2 solve the quadratic equation. Make sure x is greater than 0 as you cannot take log of a negative or 0 number.

OpenStudy (anonymous):

x=7 x=1 Didnt know that ln(x)^2 = that thannk you!

OpenStudy (ranga):

Throw away that x = 1 because ln(1) = 0 and in the original equation ln(x) appears in the denominator and therefore cannot be 0.

OpenStudy (anonymous):

Yup! Just wasted my time by checking it! Thanks

OpenStudy (ranga):

Yes. ln(A^B) = B * ln(A) So we can write 2 * ln(x) as ln(x^2)

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

Ahhhh I get it!!! Thank you!

OpenStudy (ranga):

you are welcome.